# Noncollinear Magnetic Multipoles in Collinear Altermagnets

### v1

@@ Abstract
Introduction.- Altermagnetism (AM) has recently been identified as a magnetic phase distinct from both ferromagnetism and antiferromagnetism (AF) [1, 2, 3, 3, 4, 5]. Altermagnetic order breaks time-reversal symmetry $\mathcal{T}$ by itself, but preserves the combination of $\mathcal{T}$ with some crystallographic symmetries, ensuring compensation of the net magnetic dipole moment. These symmetries are often non-symmorphic, and cannot be inversion or translation, unlike in AF. This causes otherwise spin-degenerate bands to split in most of the Brillouin zone even in the absence of spin-orbit coupling (SOC), while remaining degenerate on particular nodal lines and planes. Analogous to the nodal structure found in unconventional singlet superconductors, these spin-splitting patterns in altermagnets are labeled as d-wave, g-wave, etc., corresponding to angular momentum $l=2$, $l=4$, etc, which in turn determine the number of nodal planes in the spin-splitting throughout the Brillouin zone [6, 7, 8, 9].
@@
Collinear magnetic multipoles.- Magnetically ordered phases of crystals can be classified using magnetic space groups (MSGs), which also determine the magnetic site symmetries, i.e. the magnetic point group of each atomic site. Similar to the Wyckoff positions commonly used in crystallography, each atomic site can be assigned a magnetic Wyckoff position, which are tabulated for all MSGs [35, 36]. While the magnetic Wyckoff position tables focus on the allowed dipole moments, the magnetic site symmetries also constrain the spin density $\textbf{m}(\textbf{r})$ around an atomic site. Generally, $\textbf{m}(\textbf{r})$ can be expanded in tesseral (cubic) harmonics of order $\ell$, forming a magnetic multipole of order $\ell+1$ [37, 38]. Hence, the allowed magnetic multipoles centered on an atomic site are determined by the site symmetry as well. However, the sign and magnitude of multipoles are not determined by the site symmetry. Below, we obtain the spin density $\textbf{m}(\textbf{r})$ in real space from DFT, and calculate the magnitude of multipoles by projecting its components within atomic spheres onto different tesseral harmonics [39] (see, e.g. Ref. [40, 29, 31] for related approaches). These magnitudes are given in the atomic units of $a_{0}^{\ell}\mu_{B}$, where $a_{0}$ is the Bohr radius and $\mu_{B}$ the Bohr magneton. In the following, we use a shorthand notation and, for example, denote the octupole with a spin density pattern along $z$ that depends on position r as $m_{z}(\textbf{r})\sim xy$ as $d_{xy}m_{z}$’. The magnitude of this octupole is $\sim\int xy\cdot m_{z}(\textbf{r})\penalty 10000\ d\textbf{r}$, where the integral is over the atomic sphere (see SM for details [39]). This approach does not take into account all contributions to the total unit cell multipoles, but it allows predicting which atomic multipoles are in-phase (IP) vs. out-of-phase (OP) on different atoms within a unit cell. Thus, it circumvents the issue that unit cell multipoles do not have well-established gauge-invariant formulations due to complications in defining the position operator in periodic systems [41, 34].
Figure 1: (a) Crystal and magnetic structures of MnF2 and CrSb. The blue lines represent the screw axes, the red and green arrows represents the local $x$, $y$ coordinate axes, and the black arrows represent the spins. (b) Isosurfaces of collinear spin density ($m_{z}(\mathbf{r})$) around a magnetic ion obtained from DFT. (c) Same as (b), but for the noncollinear ($m_{x}(\mathbf{r})$ and $m_{y}(\mathbf{r})$) components, with a smaller isosurface value.
@@
Figure 2: (a) The isosurfaces for different components of DFT spin densities of MnF2. Orange and blue colors correspond to positive and negative values, and the isosurface values used for collinear ($m_{z}$) and noncollinear ($m_{x}$ and $m_{y}$) are not equal. (b) Examples of decompositon onto the tesseral harmonics of the spin density in MnF2. Red and blue data correspond to the Mn atoms on the corners and the body-center of the unit cell. Results for the rest of the multipoles are reported in the SM [39].
Spin-orbit coupling.- Both MnF2 and CrSb are collinear magnets with atomic dipoles parallel to the $c$ axis. However, any magnetic material can have a noncollinear spin density (NSD) in the general equivalent positions $(x,\penalty 10000\ y,\penalty 10000\ z)$ in the unit cell, because these positions do not have any symmetries that prohibit a local spin density $\textbf{m}(\textbf{r})$ with an arbitrary orientation. This fact, often ignored because the NSD cancels out when integrated around an atomic site in collinear magnets ($sm_{x}=\int m_{x}(\textbf{r})d\textbf{r}=sm_{y}=0$), is easy to demonstrate in DFT by simply plotting isosurfaces of the NSD. In Fig. 1c, we show the isosurfaces of $m_{x}(\textbf{r})$ and $m_{y}(\textbf{r})$ spin densities around Mn and Cr ions in MnF2 and CrSb from DFT. Even though MnF2 is a d-wave AM, it exhibits 32-poles in its real space NSD that do not cancel when summed over the two atoms, as shown in 2a. This means that even though the primary order paramater is octupolar, higher-order in-plane multipoles are dominant and affect physical observables that depend on in-plane moments. In CrSb, while non-canceling high-order multipolar behavior is also evident (see 3a), the spin density pattern is more complicated because of the partial occupation of the Cr orbitals. This necessitates projecting the NSD onto tesseral harmonics to single out the contributions from different multipoles.
@@
Figure 3: (a) The isosurfaces for different components of DFT spin densities of CrSb. Orange and blue colors correspond to positive and negative values, and the isosurface values used for collinear ($m_{z}$) and noncollinear ($m_{x}$ and $m_{y}$) are not equal. (b) Examples of decompositon onto the tesseral harmonics of the spin density in CrSb. Red and blue data correspond to the Cr atoms on the edge and the body-center of the unit cell, which are equal for IP multipoles (e.g. $d_{xy}m_{z}$) and opposite for OP multipoles (e.g. $m_{z}$ and $d_{xz}m_{x}+d_{yz}m_{y}$). Results for the rest of the multipoles are reported in the SM [39].
@@
Structural control of altermagnetism.- Finally, we discuss a third material example to elucidate how IP multipoles emerge in an antiferromagnetic-to-altermagnetic transition. The perovskite crystal structure [9, 46, 47] is an extremely flexible platform to realize antiferromagnetism and altermagnetism. In particular, the perovskite KMnF3 [46] provides a suitable playground to study the dependence of altermagnetism on structural degrees of freedom. This compound undergoes a structural transition at $\sim$88 K from cubic $Pm\bar{3}m$ to tetragonal $I4/mcm$ space group. G-type collinear magnetic order also condenses at either the same or at a close temperature [48, 49]. The magnetic order parameter transforms as the $mR_{5}^{-}$ irrep of the parent $Pm\bar{3}m.1^{\prime}$ space group. Being a zone-boundary order, it preserves anti-translation (translation followed by time reversal) symmetry and hence is an AF that does not give rise to no spin-splitting.
@@
In the tetragonal phase, the fluorine octahedra are rotated by $\sim 6^{o}$ around the $c$ axis, and the unit cell is doubled. The oxygen octahedral rotation order parameter transforms as the $R_{5}^{-}$ irrep, and has the same wavevector as the magnetic order parameter. In other words, it breaks the translational symmetry the same way that the magnetic order parameter does, and hence anti-translation is no longer a symmetry when both structural and magnetic order parameters condense. This leads to octahedral rotation induced altermagnetism in KMnF3, where the altermagnetic splitting is proportional to both the magnetic order parameter and the structural order parameter. This is evident in the DFT results. In Fig. 4, we show that the $m_{z}(\textbf{r})$ isosurfaces around each Mn ion rotate with the octahedra, and even though the total dipole density does not change much, the lowest order IP multipole, the 32-pole $g_{xy(x^{2}-y^{2})}m_{z}$, is linearly proportional to the octahedral rotation angle (Fig. 4b). This 32-pole transforms as the $m\Gamma_{1}^{+}$ irrep of $Pm\bar{3}m.1^{\prime}$, and it is induced by a trilinear coupling in the free energy: $\mathcal{F}\sim R_{5}^{-}(a,0,0)\cdot mR_{5}^{-}(a,0,0)\cdot m\Gamma_{4}^{+}$, consistent with its linear dependence on $R_{5}^{-}$ rotations [39].
@@
We note that RbMnF3 has a Néel temperature that is comparable to that of KMnF3, but it retains the cubic crystal structure at all temperatures [50]. As a result, the partially Rb-substituted compound K1-xRbxMnF3 would likely have a lower temperature structural transition that is well separated from the magnetic one, and hence could serve as a model system to study antiferromagnetic to altermagnetic phase transitions.
Tensor observables.- The symmetry of response tensors is often described using the Jahn symbol [51]. The multipoles can also be connected to a Jahn symbol [14]: for example, the magnetic octupole $r_{i}r_{j}m_{k}$ is an axial and magnetic rank-3 object symmetric under exchange of $i$ and $j$, so it is denoted by the symbol $ae[V^{2}]V$. Thus, systems with a net magnetic octupole must allow non-zero elements of $ae[V^{2}]V$ response tensors. Examples of such tensors include piezomagnetism [52, 53] and the second-order magnetoelectric effect [54, 55].
The IP multipoles in MnF2 include the collinear component $d_{xy}m_{z}$, which corresponds to the piezomagnetic tensor element $\Lambda_{xyz}$. Since it is induced by the large collinear spin components, this element remains finite even in the absence of SOC, but it is strongly suppressed in this system by Luttinger compensation, which enforces equal numbers of spin up and down electrons at zero temperature as long as the gap remains open. On the other hand, since CrSb is d-wave only in the noncollinear components, strain can induce a linear magnetization only along the $x$ and $y$ directions, and the magnitude of this effect must be proportional to $\lambda_{SOC}$. The lowest rank tensors that are linearly proportional to the collinear 32-pole $g_{yz(y^{2}-3x^{2})}m_{z}$ are rank-5 tensors with $ae[V^{2}][V^{2}]V$. These include second-order piezomagnetism (magnetization proportional to the product of two different components of strain, e.g. $\varepsilon_{yz}$ and $\varepsilon_{xx}$), and strain derivative of the second order magnetoelectric effect.
@@
Figure 4: (a) The $m_{z}(\mathbf{r})$ spin density around the Mn ions in KMnF3 is aligned with the F octahedra, and rotates without much distortion along with the octahedra. F ions are shown as grey spheres. (b) While the integrated $sm_{z}$ dipole density remains almost constant as a function of the octahedral rotation angle, the IP 32-poles $g_{xy(x^{2}-y^{2})}m_{z}$ turn on linearly with octahedral rotations. All other multipole components are reported in the SM [39].
@@
Conclusions.- Using first principles calculations, we calculated the higher order atomic multipoles in altermagnetic materials by projecting the spin density onto multipoles, which provides a simple yet robust approach to calculate these multipoles in real materials and beyond simple models. Interestingly, even in systems with seemingly spherical spin density in the collinear direction, we found that the noncollinear components provide a more elucidative illustration of the presence and relative phase of altermagnetic multipoles. Our results also showed that 32-poles, which are often ignored in the discussion of $d$-wave altermagnets, can be in fact be the dominant contribution to spin density and in certain macroscopic properties. In KMnF3, the interplay between the 32-poles and the crystal structure adds altermagnetism to the list of phenomena induced by octahedral rotations in perovskites. In summary, our work underlines that the NSD, higher order multipoles, and spin-orbit coupling are all important features of altermagnets that need to be taken into account in real materials especially when predicting their macroscopic response.

### v2

@@ Abstract
I Introduction
Altermagnetism (AM) has recently been identified as a magnetic phase distinct from both ferromagnetism and antiferromagnetism (AF) [1, 2, 3, 3, 4, 5]. Altermagnetic order breaks time-reversal symmetry $\mathcal{T}$ by itself, but preserves the combination of $\mathcal{T}$ with some crystallographic symmetries, ensuring compensation of the net magnetic dipole moment. These symmetries are often non-symmorphic, and cannot be inversion or translation, unlike in AF. This causes otherwise spin-degenerate bands to split in most of the Brillouin zone even in the absence of spin-orbit coupling (SOC), while remaining degenerate on particular nodal lines and planes. Analogous to the nodal structure found in unconventional singlet superconductors, these spin-splitting patterns in altermagnets are labeled as d-wave, g-wave, etc., corresponding to angular momentum $l=2$, $l=4$, etc, which in turn determine the number of nodal planes in the spin-splitting throughout the Brillouin zone [6, 7, 8, 9].
@@ I Introduction
II Collinear magnetic multipoles
Magnetically ordered phases of crystals can be classified using magnetic space groups (MSGs), which also determine the magnetic site symmetries, i.e. the magnetic point group of each atomic site. Similar to the Wyckoff positions commonly used in crystallography, each atomic site can be assigned a magnetic Wyckoff position, which are tabulated for all MSGs [35, 36]. While the magnetic Wyckoff position tables focus on the allowed dipole moments, the magnetic site symmetries also constrain the spin density $\textbf{m}(\textbf{r})$ around an atomic site. Generally, $\textbf{m}(\textbf{r})$ can be expanded in tesseral (cubic) harmonics of order $\ell$, forming a magnetic multipole of order $\ell+1$ [37, 38]. Hence, the allowed magnetic multipoles centered on an atomic site are determined by the site symmetry as well. However, the sign and magnitude of multipoles are not determined by the site symmetry. Below, we obtain the spin density $\textbf{m}(\textbf{r})$ in real space from DFT, and calculate the magnitude of multipoles by projecting its components within atomic spheres onto different tesseral harmonics [39] (see, e.g. Ref. [40, 29, 31] for related approaches). These magnitudes are given in the atomic units of $a_{0}^{\ell}\mu_{B}$, where $a_{0}$ is the Bohr radius and $\mu_{B}$ the Bohr magneton. In the following, we use a shorthand notation and, for example, denote the octupole with a spin density pattern along $z$ that depends on position r as $m_{z}(\textbf{r})\sim xy$ as $d_{xy}m_{z}$’. The magnitude of this octupole is $\sim\int xy\cdot m_{z}(\textbf{r})\penalty 10000\ d\textbf{r}$, where the integral is over the atomic sphere (see SM for details [39]). This approach does not take into account all contributions to the total unit cell multipoles, but it allows predicting which atomic multipoles are in-phase (IP) vs. out-of-phase (OP) on different atoms within a unit cell. Thus, it circumvents the issue that unit cell multipoles do not have well-established gauge-invariant formulations due to complications in defining the position operator in periodic systems [41, 34].
Figure 1: (a) Crystal and magnetic structures of MnF2 and CrSb. The blue lines represent the screw axes, the red and green arrows represents the local $x$, $y$ coordinate axes, and the black arrows represent the spins. (b) Isosurfaces of collinear spin density $m_{z}(\mathbf{r})$ around a magnetic ion obtained from DFT in the absence of SOC. When the SOC is introduced, the atomic dipoles do not tilt, but a nodal spin density is induced in the $m_{x}(\mathbf{r})$, $m_{y}(\mathbf{r})$ components, displayed in (c), with a smaller isosurface value than (b).
@@ II Collinear magnetic multipoles
Figure 2: (a) The isosurfaces for different components of DFT spin densities of MnF2. Orange and blue colors correspond to positive and negative values, and the isosurface values used for collinear ($m_{z}$) and noncollinear ($m_{x}$ and $m_{y}$) are not equal. (b) Examples of decompositon onto the tesseral harmonics of the spin density in MnF2. The atomic magnetic octupoles have the units of $\mu_{B}a_{0}^{2}$, while the 32-poles have the units of $\mu_{B}a_{0}^{4}$, where $\mu_{B}$ is the Bohr magneton and $a_{0}$ is the Bohr radius. Red and blue data correspond to the Mn atoms on the corners and the body-center of the unit cell, which are equal for IP multipoles (e.g. $d_{xy}m_{z}$) and opposite for OP multipoles (e.g. $m_{z}$ and $d_{xz}m_{x}+d_{yz}m_{y}$). Results for the rest of the multipoles are reported in the SM [39].
III Spin-orbit coupling
Both MnF2 and CrSb are collinear magnets with atomic dipoles parallel to the $c$ axis. However, any magnetic material can have a noncollinear spin density (NSD) in the general equivalent positions $(x,\penalty 10000\ y,\penalty 10000\ z)$ in the unit cell, because these positions do not have any symmetries that prohibit a local spin density $\textbf{m}(\textbf{r})$ with an arbitrary orientation. This fact, often ignored because the NSD cancels out when integrated around an atomic site in collinear magnets ($sm_{x}=\int m_{x}(\textbf{r})d\textbf{r}=sm_{y}=0$), is easy to demonstrate in DFT by simply plotting isosurfaces of the NSD. In Fig. 1c, we show the isosurfaces of $m_{x}(\textbf{r})$ and $m_{y}(\textbf{r})$ spin densities around Mn and Cr ions in MnF2 and CrSb from DFT. Even though MnF2 is a d-wave AM, it exhibits 32-poles in its real space NSD that do not cancel when summed over the two atoms, as shown in 2a. This means that even though the primary order paramater is octupolar, higher-order in-plane multipoles are dominant and affect physical observables that depend on in-plane moments. In CrSb, while non-canceling high-order multipolar behavior is also evident (see 3a), the spin density pattern is more complicated because of the partial occupation of the Cr orbitals. This necessitates projecting the NSD onto tesseral harmonics to single out the contributions from different multipoles.
@@ III Spin-orbit coupling
Figure 3: (a) The isosurfaces for different components of DFT spin densities of CrSb. Orange and blue colors correspond to positive and negative values, and the isosurface values used for collinear ($m_{z}$) and noncollinear ($m_{x}$ and $m_{y}$) are not equal. (b) Examples of decompositon onto the tesseral harmonics of the spin density in CrSb. The atomic magnetic octupoles have the units of $\mu_{B}a_{0}^{2}$, while the 32-poles have the units of $\mu_{B}a_{0}^{4}$, where $\mu_{B}$ is the Bohr magneton and $a_{0}$ is the Bohr radius. Red and blue data correspond to the Cr atoms on the edge and the body-center of the unit cell, which are equal for IP multipoles (e.g. $d_{xy}m_{x}+d_{x^{2}-y^{2}}m_{y}$) and opposite for OP multipoles (e.g. $d_{z^{2}}m_{z}$). Results for the rest of the multipoles are reported in the SM [39].
IV Structural control of altermagnetism
@@
Finally, we discuss a third material example to elucidate how IP multipoles emerge in an antiferromagnetic-to-altermagnetic transition. The perovskite crystal structure [9, 46, 47] is an extremely flexible platform to realize antiferromagnetism and altermagnetism. In particular, the perovskite KMnF3 [46] provides a suitable playground to study the dependence of altermagnetism on structural degrees of freedom. This compound undergoes a structural transition at $\sim$88 K from cubic $Pm\bar{3}m$ to tetragonal $I4/mcm$ space group. G-type collinear magnetic order also condenses at either the same or at a close temperature [48, 49]. The magnetic order parameter transforms as the $mR_{5}^{-}$ irrep of the parent $Pm\bar{3}m.1^{\prime}$ space group. Being a zone-boundary order, it preserves anti-translation (translation followed by time reversal) symmetry and hence is an AF that does not give rise to no spin-splitting.
@@
In the tetragonal phase, the fluorine octahedra are rotated by $\sim 6^{o}$ around the $c$ axis, and the unit cell is doubled. The oxygen octahedral rotation order parameter transforms as the $R_{5}^{-}$ irrep, and has the same wavevector as the magnetic order parameter. In other words, it breaks the translational symmetry the same way that the magnetic order parameter does, and hence anti-translation is no longer a symmetry when both structural and magnetic order parameters condense. This leads to octahedral rotation induced altermagnetism in KMnF3, where the altermagnetic splitting is proportional to both the magnetic order parameter and the structural order parameter. This is evident in the DFT results. In Fig. 4, we show that the $m_{z}(\textbf{r})$ isosurfaces around each Mn ion rotate with the octahedra, and even though the total dipole density does not change much, the lowest order IP multipole, the 32-pole $g_{xy(x^{2}-y^{2})}m_{z}$, is linearly proportional to the octahedral rotation angle (Fig. 4b). This 32-pole transforms as the $m\Gamma_{1}^{+}$ irrep of $Pm\bar{3}m.1^{\prime}$, and it is induced by a trilinear coupling in the free energy: $\mathcal{F}\sim R_{5}^{-}(a,0,0)\cdot mR_{5}^{-}(a,0,0)\cdot m\Gamma_{4}^{+}$, consistent with its linear dependence on $R_{5}^{-}$ rotations [39]. A similar rotation-induced altermagnetic phase with the same rotation mode (albeit with a rhombohedral structure) has also been studied in BiFeO3 [50], attesting to the fact that it is a common feature in many magnetic perovskites.
@@
We note that the related compound RbMnF3 has a Néel temperature that is comparable to that of KMnF3, but it retains the cubic crystal structure at all temperatures [51]. As a result, the partially Rb-substituted compound K1-xRbxMnF3 would likely have a lower temperature structural transition that is well separated from the magnetic one, and hence could serve as a model system to study antiferromagnetic to altermagnetic phase transitions.
@@
Figure 4: (a) The $m_{z}(\mathbf{r})$ spin density around the Mn ions in KMnF3 is aligned with the F octahedra, and rotates without much distortion along with the octahedra. F ions are shown as grey spheres. (b) While the integrated $sm_{z}$ atomic dipole density remains almost constant as a function of the octahedral rotation angle, the IP 32-poles $g_{xy(x^{2}-y^{2})}m_{z}$ turn on linearly with octahedral rotations. All other multipole components are reported in the SM [39].
V Tensor observables
The symmetry of response tensors is often described using the Jahn symbol [52]. The multipoles can also be connected to a Jahn symbol [14, 53]: for example, the magnetic octupole $r_{i}r_{j}m_{k}$ is an axial and magnetic rank-3 object symmetric under exchange of $i$ and $j$, so it is denoted by the symbol $ae[V^{2}]V$. Thus, systems with a net magnetic octupole must allow non-zero elements of $ae[V^{2}]V$ response tensors. Examples of such tensors include piezomagnetism [54, 55] and the second-order magnetoelectric effect [56, 57].
The IP multipoles in MnF2 include the collinear component $d_{xy}m_{z}$, which corresponds to the piezomagnetic tensor element $\Lambda_{xyz}$, which we define through the linear relationship between the applied strain $\epsilon_{ij}$ and the induced magnetic moment $M_{k}$ as
$M_{k}=\Lambda_{ijk}\epsilon_{ij}.$ (1)
Since $\Lambda_{xyz}$ is induced by the large collinear spin components, this element remains finite even in the absence of SOC, but it is strongly suppressed in this system by Luttinger compensation, which enforces equal numbers of spin up and down electrons at zero temperature as long as the gap remains open [3, 58, 59]. In order to make this prediction more evident, in Fig. 5(a) we plot the magnitude of $\Lambda_{xyz}$ as a function of the multipole $d_{xy}m_{z}$. We calculated the piezomagnetic coefficients by straining the unit cell in our DFT calculations, relaxing the atomic positions under fixed strain, and performing a self-consistent electronic structure calculation to obtain the value of magnetization. In the calculation without SOC, the $d_{xy}m_{z}$ multipole is nonzero but the $\Lambda_{xyz}$ piezomagnetic coefficient is zero within the scale of this plot, as expected. This, however, is not the case for other components of $\Lambda$. For instance, the $\Lambda_{xzy}$ component, which is not symmetry allowed without SOC depends on the corresponding magnetic octupole linearly for small values, with only some deviation from the linear behavior for the actual value of $\lambda_{SOC}$. Note that this component is not affected from the compensation either, because the magnetization it leads to is formed by the tilting of local atomic moments.
On the other hand, since CrSb is d-wave only in the noncollinear components, strain can induce a linear magnetization only along the $x$ and $y$ directions, and the magnitude of this effect must be proportional to $\lambda_{SOC}$. The lowest rank tensors that are linearly proportional to the collinear 32-pole $g_{yz(y^{2}-3x^{2})}m_{z}$ are rank-5 tensors with $ae[V^{2}][V^{2}]V$. These include second-order piezomagnetism (magnetization proportional to the product of two different components of strain, e.g. $\varepsilon_{yz}$ and $\varepsilon_{xx}$), and strain derivative of the second order magnetoelectric effect.
Figure 5: The dependence of the piezomagnetic tensor components to the corresponding magnetic octupoles in MnF2. Different data points are generated by repeating the DFT calculations for different SOC strengths. (a) $\Lambda_{xyz}$, which relates the $z$ component of magnetization to xyxy shear strain, is symmetry allowed even in the absence of SOC, however, it is forced to be zero by Luttinger compensation at zero temperature since there is a gap. However, the corresponding atomic octupole $d_{xy}M_{z}$ is not affected by the compensation and is nonzero for all values of SOC. The interplay between the compensation and response leads to a nontrivial dependence of this component of the piezomagnetic tensor on the atomic multipole strength. (b) For small values of SOC, the $\Lambda_{xzy}$ component, which relates the $y$ component of magnetization to xzxz shear strain, linearly depends on the $d_{xz}M_{y}$ magnetic octupole, and both quantities are zero in the absence of SOC.
VI Conclusions
@@
Using first principles calculations, we calculated the higher order atomic multipoles in altermagnetic materials by projecting the spin density onto multipoles, which provides a simple yet robust approach to calculate these multipoles in real materials and beyond simple models. Interestingly, even in systems with seemingly spherical spin density in the collinear direction, we found that the noncollinear components provide a more elucidative illustration of the presence and relative phase of altermagnetic multipoles. Our results also showed that 32-poles, which are often ignored in the discussion of $d$-wave altermagnets, can be in fact be the dominant contribution to spin density and in certain macroscopic properties. In KMnF3, the interplay between the 32-poles and the crystal structure adds altermagnetism to the list of phenomena induced by octahedral rotations in perovskites. In summary, our work underlines that the NSD, higher order multipoles, and spin-orbit coupling are all important features of altermagnets that need to be taken into account in real materials especially when predicting their macroscopic response.
VII Acknowledgments